# Understanding entropy production via a thermal zero-player game

### v2

@@ Abstract
A new thermal bath scheme for Ising-Conway Entropy Game (ICEg) is introduced. New game moves in sampling the given temperature is achieved via Monte Carlo dynamics of both Metropolis and Glauber as a stochastic game. This kind of approach makes the game an ideal tool for demonstrating thermal dependency of entropy production in a novel way. Using this new approach, Ising-Conway Entropy game’s rate of entropy production depending on different temperatures are explored. Thermalized game is shown to be physically interesting and plausible test bed for studying complex dynamical systems in classical statistical mechanics, that is conceptually simple, pedagogically accessible, yet realistic.
I Introduction
The second law of thermodynamics dictates systems evolve from ordered to disordered states, preferring maximal entropy for reversible systems and cellular automata provides a test bed for this kind of behavior without any physical system in mind [1, 2, 3]. One of the inception of cellular automa, Conway’s game is the first [4, 5, 6, 7, 8, 9] game with simple rules that can generate discrete complex patterns over dynamical modes. In this direction, a recent Ising-Conway Entropy Game ICEg [10] introduced a zero-player game for statistical physics where by Ising model like game [10] shows entropy increase, starting from a corner cells (occupied spins) in a manner reminiscent of Conway’s game and Ising model [11], albeit simpler rules. Temperature for Conway’s game is introduced earlier [12], however it only introduce stochasticity in Conway’s rules, i.e., a basic noise term. On the other hand, introducing thermal bath in molecular dynamics is widely studied in the literature [13], Approaches usually modifies dynamical equations appropriately introduce temperature. The zero player game like ICEg using Monte Carlo [14, 15] approach in accepting moves depending on temperature is suitable. Using Metropolis [16] and Glauber [17, 18] spin-flip dynamics [14, 15] is possible. This novel approach goes beyond to be simple noise term only but induce a dynamics that samples the given Potential Energy Surface (PES) of the corresponding temperature of the system in heat bath. This novel approach enables us to study a game within statistical physics context [19]. entropy production is also studied in this context [20], though in the thermal zero-player game, payoffs are less obvious and manifests in the rules.
@@ I Introduction
II Thermal Zero Player Game
@@
Example evolution of game for both Metropolis and Glauber dynamics is given by Figure 1 and Figure 2, at $\beta=0.9$, $N=150$ and $M=45$, evolution of lattice sites over time is given by stacked up fashion. We observe how diffusion behavior manifests over time. All results should be averaged over repeated game plays. In the results we use 100 repeats. Note that, from game theoretic perspective payoffs appears as acceptable move leading to configuration that satisfiy the given temperature.
III Entropy for the game
Entropy is an elusive concept and it manifest in different settings, originating from Thermodynamics of heat engines [22]. Shannon’s information entropy is specially suitable for discrete dynamics and covers configurational entropy from physics perspective [23], compare to Gibbs-Boltzmann entropies [1]. Quantum [24] and black-hole entropies [25] manifests differently. In our context, we take the largest extent of the lattice dynamics we presented above as a surrogate to Entropy of the system. We identify this extent as $S(t)$ on the given lattice $L(t)$ of size $N$ as described above
@@ III Entropy for the game
IV Transition to entropic regime
Evolution of entropy measure for different temperatures $\beta=0.01,0.1,0.2,0.4,0.6,0.8,0.9$ are produced with two different lattice settings of $(N,M)$, (100, 20) and (150,45). In all cases, clear transition to entropic regime in all temperatures are observed for both Metropolis and Glauber dynamics.
@@ IV Transition to entropic regime
V Entropy production
@@
On Figure 7, entropy production over different temperatures are given, as expected production increases up to a critical temperature and then saturates. $S_{prod}$ consistently higher in the case of Metropolis dynamics. This is due to fact that Glauber dynamics differentiates entropy dynamics more pronounced with increasing temperature.
VI Conclusions
In this work we have introduced thermal bath for Ising-Conway Entropy game with a novel coupling scheme. As a pedagogically quite accessible zero-player game for statistical mechanics, playing the game with different settings provides quite valuable insights conceptually for entropy in classical statistical physics. The rate of increase in entropy with different temperatures are shown to be more distict for Glauber dynamics compare to Metropolis dynamics, this indicates, using Glauber dynamics should be preferred in general. Observation that entropy production is bound to have a critical temperature that after it doesn’t change. This insight is shown for the first time with such a simple system without resorting to free-energy computations.

### v3

@@ Abstract
A new thermal bath scheme for Ising-Conway Entropy Game (ICEg) is introduced. New game moves for sampling the given temperature are achieved via Monte Carlo dynamics of both Metropolis and Glauber in a stochastic game. This kind of approach makes the game an ideal tool for demonstrating thermal dependency of entropy production in a novel way. Using this new approach, Ising-Conway Entropy game’s rate of entropy production depending on different temperatures is explored. Thermalized game is shown to be physically interesting and a plausible test bed for studying complex dynamical systems in classical statistical mechanics, that is conceptually simple, pedagogically accessible, yet realistic.
  • June 2025
1 Introduction
The second law of thermodynamics dictates systems evolve from ordered to disordered states, preferring maximal entropy for reversible systems and cellular automata provide a test bed for this kind of behavior without any physical system in mind [13, 18, 32]. One of the inceptions of cellular automa, Conway’s game is the first [8, 7, 10, 11, 9, 12] game with simple rules that can generate discrete complex patterns over dynamical modes. In this direction, a recent Ising-Conway Entropy Game ICEg [29] introduced a zero-player game for statistical physics where by Ising model like game [29] shows entropy increase, starting from a corner cells (occupied spins) in a manner reminiscent of Conway’s game and Ising model [16], albeit simpler rules. Temperature for Conway’s game is introduced earlier [26], however it only introduces stochasticity in Conway’s rules, i.e., a basic noise term. On the other hand, introducing thermal bath in molecular dynamics is widely studied in the literature [15], Approaches usually modify dynamical equations to appropriately introduce temperature. The zero player game like ICEg using Monte Carlo [3, 23] approach in accepting moves depending on temperature is suitable. Using Metropolis [22] and Glauber [14, 28] spin-flip dynamics [3, 23] is possible. This novel approach goes beyond to be simple noise term only but induces a dynamics that samples the given Potential Energy Surface (PES) of the corresponding temperature of the system in heat bath. This novel approach enables us to study a game within statistical physics context [1]. Entropy production is also studied in this context [6], though in the thermal zero-player game, payoffs are less obvious and manifest in the rules.
@@ 1 Introduction
2 Thermal Zero Player Game
@@
Example evolution of game for both Metropolis and Glauber dynamics is given by Figure 1 and Figure 2, at $\beta=0.9$, $N=150$ and $M=45$, evolution of lattice sites over time is given by stacked up fashion. We observe how diffusion behavior manifests over time. All results should be averaged over repeated game plays. In the results we use 100 repeats. Note that, from game theoretic perspective payoffs appear as acceptable moves leading to configuration that satisfies the given temperature.
3 Entropy for the game
Entropy is an elusive concept and it manifests in different settings, originating from Thermodynamics of heat engines [4]. Shannon’s information entropy is specially suitable for discrete dynamics and covers configurational entropy from physics perspective [27], compare to Gibbs-Boltzmann entropies [13]. Quantum [31] and black-hole entropies [2] manifest differently. In our context, we take the largest extent of the lattice dynamics we presented above as a surrogate to Entropy of the system. We identify this extent as $S(t)$ on the given lattice $L(t)$ of size $N$ as described above
@@ 3 Entropy for the game
4 Transition to entropic regime
Evolution of entropy measure for different temperatures $\beta=0.01,0.1,0.2,0.4,0.6,0.8,0.9$ are produced with two different lattice settings of $(N,M)$, (100, 20) and (150,45). In all cases, clear transition to entropic regime in all temperatures is observed for both Metropolis and Glauber dynamics.
5 Entropy production
@@
On Figure 7, entropy production over different temperatures are given, as expected production increases up to a critical temperature and then saturates. $S_{prod}$ consistently higher in the case of Metropolis dynamics. This is due to the fact that Glauber dynamics differentiates entropy dynamics more pronouncedly with increasing temperature.
6 Conclusions
In this work we have introduced a thermal bath for the Ising-Conway Entropy game with a novel coupling scheme. As a pedagogically quite accessible zero-player game for statistical mechanics, playing the game with different settings provides quite valuable insights conceptually for entropy in classical statistical physics. The rate of increase in entropy with different temperatures is shown to be more distinct for Glauber dynamics compared to Metropolis dynamics. This observation indicates that using Glauber dynamics should be preferred in general. Entropy production has a critical temperature after which it doesn’t change. This insight is shown for the first time with such a simple system without resorting to free-energy computations.